two birds fly in the same direction if one goes twice as fast as the other and it takes them 3 hours to be 57 miles apart find their speeds
Answer by TimothyLamb(4379)

x = speed fast bird
y = speed slow bird
---
s = d/t
---
relative speed of birds:
x = 2y
x - y = 57/3
2y - y = 57/3
y = 57/3
x = 2*57/3
---
answer:
x = speed fast bird = 38 mph
y = speed slow bird = 19 mph
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php